write each quadratic equation in standard form.
1. y=3(x+2)squared-10 2.y= -(x+4)squared-1 3.y=(x-5)squared+2
my question is how do i solve these
sorry about the "squared" i didnt know how to put the 2 up in the corner-- in squared form..

Answer by ankor@dixie-net.com(22740)

1. y = 3(x+2)^2 - 10
FOIL (x+2)(x+2)
y = 3(x^2 + 2x + 2x + 4) - 10
y = 3(x^2 + 4x + 4) - 10
y = 3x^2 + 12x + 12 - 10
y = 3x^2 + 12x + 2
:
:
2.y = -(x+4)^2 - 1
FOIL(x+4)(x+4)
y = -1(x^2 + 4x + 4x + 16) - 1
y = -1(x^2 + 8x + 16) - 1
y = -x^2 - 8x - 16 - 1
y = -x^2 - 8x - 17
:
3.y = (x-5)^2 + 2
You should be able to do this one, just like we did the 1st two